OpenStudy (anonymous):

Find f g h f(x)=x+1 g(x)=2x h(x)=x-1

OpenStudy (anonymous):

are you trying to find \[f(g(h(x)))\]?

OpenStudy (anonymous):

try you

OpenStudy (anonymous):

start with g(h(x)) you know that g(x) =2x so replace every x in the g function with h(x)=x-1

OpenStudy (anonymous):

f(g(x-1)))

OpenStudy (anonymous):

then

OpenStudy (anonymous):

f(x)= x+1 fg(x)= (2x)+1 fgh(x)=2(x-1)+1

OpenStudy (anonymous):

i thing is 2x-1

OpenStudy (anonymous):

for f(g(x-1)) you need to plug (x-1) in for x in the g(x) function. g(x)=2x whereas g(x-2)= 2(x-2)=2x-4 Now if you do f(g(x-2)) that is equal to f(2x-4) we know that f(x)=x+1 therefore f(2x-4)=(2x-4) + 1 = 2x -3 Follow the same steps with your problem

OpenStudy (anonymous):

you are right 2(x-1)+1 2x-2+1 2x-1